LBNL Masthead A-Z IndexBerkeley Lab mastheadU.S. Department of Energy logoPhone BookJobsSearch
2014 Shutdown: Week Three Print

We're in our third week of the shutdown at the ALS and are making very good progress on all the prioritized tasks that we need to accomplish by early July. The old storage ring radio frequency crowbar system has been removed, and the new high-voltage switch is being built in-situ. At the current pace, high-voltage testing of the new system will begin in early June.

In other work, the replacement of the  beam position monitor buttons is nearly complete, and the installation of the new vortex-based water flow sensors is 50% complete. The storage ring realignment is just finishing, though much survey and alignment work remains on checking and aligning top-off Injection safety apertures and moving beamline front-ends into optimal positions as many photon source points have been moved during the process..

The transformer in the QFA power supply that caused the unexpected downtime prior to the shutdown is currently being repaired and is on track to be here and installed by the end of the shutdown. There are many ongoing accelerator improvement activities as well as at least six beamlines that are undergoing optical system upgrades and maintenance. We will update these activities next month.